OpenStudy (anonymous):

what is the remainder of the division problem shown below (x^3+2x^2-6x-9)/(x+2)

OpenStudy (amistre64):

would you agree that: \[\frac{p(x)}{x-a}=q(x)+\frac{r(x)}{x-a}\]

OpenStudy (amistre64):

for example \[\frac 54=1+\frac14\]

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then lets simplify it, by multiply both sides by x-a \[p(x)=(x-a)~q(x)+r(x)\] do we still agree?

OpenStudy (anonymous):

yes we do

OpenStudy (amistre64):

now, let x=a \[p(a)=\underbrace{(a-a)}_{\color{red}{=0}}~q(a)+r(a)\] \[p(a)=r(a)\]

OpenStudy (amistre64):

now, how can we apply this to your question

OpenStudy (amistre64):

(x^3+2x^2-6x-9) / (x+2) = q(x) + r(x)/(x+2) x^3+2x^2-6x-9 = (x+2) q(x) + r(x) when does x+2 = 0?

OpenStudy (anonymous):

ok thanks i got it from here

OpenStudy (amistre64):

good luck :)